OpenStudy (anonymous):

describe how you would graph the line y=-3/2x+4 using the slope-intercept method.

OpenStudy (anonymous):

well this equation belongs to this form y = mx + b where m = -3/2 b = 4

OpenStudy (anonymous):

to graph we give X adecuate values, for example: I know that "b" is the intercept between graph and y-axis, then x = 0 a first pint is ( 0 ; 4)I

OpenStudy (anonymous):

to get a second point we give x = 2 then y = (-3/2) 2 + 4 = -3 + 4 y = 1 so we have 2nd point (2 ; 1) finally we make the graph

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

well it was a pleasure to help you